A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?
1
Expert's answer
2020-07-22T17:55:55-0400

-7x+3y=-21.5

3y=7x-21.5

y=7/3x-7.5

the central street passes through the origin

y=ax

straight lines are perpendicular if 7/3a=-1

a=-3/7

 the equation of the central street PQ y=-3/7x